OpenStudy (anonymous):

A grocery store sells a mixture of peanuts and raisins for $1.75 per pound. if peanuts cost $1.25 per pound and raisins cost $2.75 per pound, what amount of raisins and peanuts go into 1 pound of the mixture?

OpenStudy (anonymous):

Let R be raisins quantity and P be peanuts quantity. R + P = 1 1.25*P + 2.75*R = 1.75 Solve for R and P.

OpenStudy (anonymous):

please just answer the question, i will give you a medal i'm tired and need some sleep, please just help me out here

OpenStudy (anonymous):

1.25*(1-R) + 2.75*R = 1.75 1.25 + 1.5R = 1.75 1.5R = 1.75 - 1.25 R = 0.5/1.5 R = 0.33 P = 0.67

OpenStudy (anonymous):

Medal?